# MedSegMamba: 3D CNN-Mamba Hybrid Architecture for Brain Segmentation

### v2

@@ 3 Results
While both Mamba-based models showed competitive performance, MedSegMamba demonstrated slight improvements over SegMambaBot, and has approximately 20% fewer parameters while requiring similar GPU memory during inference (1.988 GiB for MedSegMamba vs. 1.932 GiB for SegMambaBot).
As shown in Figure 8, MedSegMamba demonstrates superior performance while remaining parameter efficient compared to the other 3D patch-based methods. While FastSurfer’s fully convolutional architecture requires very few parameters, each of its three sub-models requires more than double the GPU memory during inference (just under 5 GiB) compared to the 3D models, all of which operate with under 2 GiB of GPU memory.
Figure 7: 2D slices of a sample segmented by each method.
Figure 8: Plot of millions of parameters versus overall dice score for each model.
@@ 4 Discussion
The transition from a Transformer bottleneck to a Mamba-based bottleneck enhances memory efficiency, enabling the use of larger convolutional layers on identical hardware. This enhancement boosts local feature extraction capabilities while preserving global context extraction, resulting in both SegMambaBot and MedSegMamba outperforming TABSurfer. Comparing SegMambaBot and MedSegMamba, the SS3D module allows MedSegMamba to attain comparable overall performance in terms of DSC and VS metrics but offers superior region boundary delineation as reflected by the ASSD, all while utilizing significantly fewer parameters than the Tri-oriented Mamba module. Increasing the number of selective scan operations to process variously unraveled sequences can achieve performance parity with larger modules that process fewer sequences.
Compared to FastSurfer’s standard 2.5D approach, our 3D inputs preserve more intricate spatial relationships within the continuity of anatomy compared to 2D slices. The 3D strided reconstruction process, which employs voting on the predicted class, further mitigates noise artifacts present in the ground truth. This improvement is observed qualitatively when comparing the noisier segmentations of FastSurfer to the smoother outputs of the 3D patch-based models.
@@ 4.1 Limitations
The primary limitation of this approach is that MedSegMamba exhibits slower performance than SegMambaBot, despite its lower parameter count. However, this issue can be mitigated by increasing the step size during the strided reconstruction process from 16 to 32, resulting in a roughly 4x speedup, reducing the processing time per scan to around 22 seconds with minimal impact on performance. When both using a step size of 32, MedSegMamba is only a few seconds slower than SegMambaBot per scan, making them roughly comparable in terms of overall efficiency while retaining the superior performance afforded by the SS3D module.
@@ 4.2 Conclusion
These results showcase the advantages of our hybrid architecture combined with a 3D patch-based approach. Future research should explore the application of our SS3D module across different architectures, including its integration at various latent space levels throughout the encoder, as seen in SegMamba’s original design. Ablation studies focused on determining the optimal number of scanning directions or unraveling methods for 3D segmentation may further enhance our model’s performance.

### v3

@@ 3.1 Subcortical Segmentation
Figure 5: 3D renderings of a sample’s subcortical segmentation by each method.
Figure 6: (a) Plot of millions of parameters versus subcortical segmentation dice score for each model. (b) Plot of millions of parameters versus hippocampal subfield segmentation dice score for each model.
@@
While both Mamba-based models showed competitive performance, MedSegMamba demonstrated slight improvements over SegMambaBot and has approximately 20% fewer parameters while requiring similar GPU memory during inference (1.988 GiB for MedSegMamba vs. 1.932 GiB for SegMambaBot).
As shown in Figure 6a, MedSegMamba demonstrates superior performance while remaining parameter efficient compared to the other 3D patch-based methods. While FastSurfer’s fully convolutional architecture requires very few parameters, each of its three sub-models requires more than double the GPU memory during inference (just under 5 GiB) compared to the 3D models, all of which operate with under 2 GiB of GPU memory.
3.2 Hippocampal Subfield Segmentation
Figure 7: 3D renderings of a sample’s hippocampal subfield segmentation by each method.
Figure 8: 2D slices of a sample’s hippocampal subfield segmentation by each method are shown. Large regions in the right hippocampus are missing from TABSurfer’s segmentation, as indicated by the red arrow in the axial view. In SegMambaBot’s segmentation, the left hippocampal amygdala transition area is missing, as shown in the axial view, and the right presubiculum is slightly undersegmented, as shown in the sagittal right view.
Dataset Model DSC ↑ VS ↑ ASSD* ↓
ADNI MedSegMamba 0.80415±0.047 0.96207±0.014 0.22819±0.068
SegMambaBot 0.76976±0.045 0.92309±0.014
TABSurfer 0.61932±0.037 0.75894±0.011
AIBL MedSegMamba 0.80935±0.037 0.96511±0.012 0.21981±0.055
SegMambaBot 0.77561±0.032 0.92525±0.010
TABSurfer 0.62564±0.031 0.76095±0.009
BGSP MedSegMamba 0.81948±0.023 0.96375±0.010 0.21061±0.037
SegMambaBot 0.78562±0.018 0.92611±0.008
TABSurfer 0.63435±0.016 0.76336±0.009
IXI MedSegMamba 0.80858±0.035 0.96461±0.010 0.22449±0.052
SegMambaBot 0.77445±0.036 0.92566±0.010
TABSurfer 0.62394±0.033 0.75991±0.009
NIFD MedSegMamba 0.81438±0.045 0.96871±0.008 0.22214±0.062
SegMambaBot 0.77286±0.051 0.92698±0.007
TABSurfer 0.61505±0.043 0.76174±0.006
OAS1 MedSegMamba 0.79368±0.077 0.95663±0.027 0.27493±0.288
SegMambaBot 0.75796±0.072 0.91743±0.027
TABSurfer 0.61299±0.061 0.75517±0.020
Overall MedSegMamba 0.80683±0.046 0.96299±0.015 0.22960±0.115
SegMambaBot 0.77219±0.043 0.92379±0.014
TABSurfer 0.62276±0.037 0.75979±0.012
Table 2: Comparing MedSegMamba, SegMambaBot, and TABSurfer metrics across the hippocampal subfield segmentation datasets. Bold text indicates superior performance. ↑ indicates that higher numbers correspond to better performance and ↓ indicates that lower numbers correspond to better performance. *ASSD is not shown for SegMambaBot and TABSurfer because it cannot be computed for the classes which SegMambaBot and TABSurfer failed to learn.
Metrics evaluating MedSegMamba, SegMambaBot, and TABSurfer’s hippocampal segmentations are shown in Table 2. MedSegMamba’s improvements over both SegMambaBot and TABSurfer are more notable for this task than for subcortical segmentation, as seen in Figure 6b. Likely due to the smaller structures present in this task, MedSegMamba was the only network that learned to predict all 25 classes, while SegMambabot only learned 24 classes, and TABSurfer converged on predicting just 20 classes. This resulted in significant improvements over both models in DSC and VS. ASSD cannot be computed for missing classes, so statistical tests are not applicable between the models’ overall ASSD. However, even when ignoring the missing classes, SegMambaBot and TABSurfer get an average ASSD of 0.23380 and 0.25512, respectively, which are both still inferior to MedSegMamba’s average of 0.22960. The 3D and 2D visualizations of a sample’s hippocampus segmentation by each method are visualized in Figure 7 and Figure 8. TABSurfer failed to segment numerous classes, leaving large regions missing. While SegMambaBot and MedSegMamba appear similar, SegMambaBot is missing the left hippocampal amygdala transition area and slightly under-segments some regions, leaving MedSegMamba as the only model to succeed in segmenting each subregion.
@@ 4 Discussion
The transition from a Transformer bottleneck to a Mamba-based bottleneck significantly enhances memory efficiency, thereby enabling the use of more convolutional layers for the encoder and decoder on identical hardware. This enhancement boosts local feature extraction capabilities while preserving global context extraction, resulting in both SegMambaBot and MedSegMamba outperforming TABSurfer. Comparing SegMambaBot and MedSegMamba in subcortical segmentation, the SS3D module allows MedSegMamba to attain comparable overall performance in terms of DSC and VS metrics but offers superior region boundary delineation as reflected by the ASSD, all while utilizing significantly fewer parameters than the Tri-oriented Mamba module. Increasing the number of selective scan operations to process variously unraveled sequences can achieve performance parity with larger modules that process fewer sequences.
Compared to FastSurfer’s standard 2.5D approach, our utilization of 3D inputs preserves more intricate spatial relationships within the anatomy’s continuity than 2D slices. The 3D strided reconstruction process, which aggregates model outputs from shifted overlapping patches, further mitigates noise artifacts present in the ground truth. This improvement is observed qualitatively when comparing the noisier segmentations of FastSurfer to the smoother outputs of the 3D patch-based models.
The hippocampal subfield segmentation experiment also provides evidence of the SS3D module’s ability to process finer details compared to the Tri-oriented Mamba module. While SegMambaBot and TABSurfer failed this task, MedSegMamba was able to reliably segment all the small and intricate subregions.
@@ 4.1 Limitations
The primary limitation of this approach is that MedSegMamba exhibits slower performance than SegMambaBot despite its lower parameter count, most notably during the subcortical segmentation pipeline. However, this issue can be mitigated by increasing the step size during the strided reconstruction process from 16 to 32, resulting in a roughly 4x speedup, reducing the processing time to around 22 seconds per scan with minimal impact on performance. When both use a step size of 32, MedSegMamba’s subcortical segmentation is only a few seconds slower than SegMambaBot per scan, making them roughly comparable in terms of overall efficiency while retaining the superior performance afforded by the SS3D module.
@@ 4.2 Conclusion
These results showcase the advantages of this novel hybrid architecture combined with 3D patch-based processing. Future research should explore the application of our SS3D module across different architectures, including its integration at various latent space levels throughout the encoder, as seen in SegMamba’s original design. Ablation studies focused on determining the optimal number of scanning directions or unraveling methods for 3D segmentation may further enhance our model’s performance.